Exterminator:

Subtract. Write your answer in simplest form 5/4-8/10= plz help Thx!

snowflake0531:

\[\frac{ 5 }{ 4 } - \frac{ 8 }{ 10 }\] To subtract, we first have to make the denominator equal So, can you first find the LCM of 4 and 10?

Exterminator:

40?

snowflake0531:

There's actually a number smaller than that

Exterminator:

ah 20

snowflake0531:

Yep, and so we have to get both fractions' denominator to 20, can you do that?

Exterminator:

5/20-8/10

Exterminator:

this?

snowflake0531:

You cannot only change the denominator, one sec

snowflake0531:

\[\frac{ 5 }{ 4 } * \frac{ 5 }{ 5 } - \frac{ 8 }{ 10 } * \frac{ 2 }{ 2 }\] can you simplify that?

Exterminator:

nvm i figured it out thx!!!!!!!!!!!!!!!!!!!

snowflake0531:

yw~!!